# Enhancing Underwater Imaging with 4-D Light Fields: Dataset and Method

### v1

@@ V-C Network Architecture Evaluation
The scores of PSNR and SSIM metrics of different methods are reported in Table II. Our proposed method demonstrates the best performance as presented. Unlike 2-D RGB-based methods, the geometric information from the 4-D LFs proves advantageous in improving underwater imaging quality. Our specific strategy enables our framework to outperform two other LF-based methods, specifically due to our meticulous design for underwater environments. As for reference time, LF methods, including DistgSSR, MSPNet, and ours, have the capability to enhance all views of a scene simultaneously, whereas other methods require processing each image individually.
@@ V-D Ablation Study
TABLE IV: Quantitative results of ablation study about different numbers of iterative stages. Numbers 1 to 5 represent the iterative stages
@@
2) Evaluation of Depth Estimation. Severe underwater degradation can compromise the content of images, leading to poor depth maps. As illustrated in the first and third columns of Fig. 12, the depth maps derived from estimated underwater images are subpar and lack a clear structural hierarchy. Conversely, as shown in the second and fourth columns of Fig. 12, the depth maps estimated from GT and restored image demonstrate superior results, providing sharp edges and preserving smoothness in areas of uniform depth. These results highlight the effectiveness of our framework with OccUnNet in providing better depth maps of enhanced LF images.
@@
We further conduct experiments on our proposed network with different stages (1, 2, 3, 4, 5). We train our model on our dataset with 5 × 5 angular resolution. Table IV shows the results of different stages. Overall, adding more processing stages can enhance performance. However, beyond four stages, the improvement in enhancement performance becomes less significant. Additionally, more stages mean more parameters, which can slow down the model running. Therefore, we have made a trade-off between the model’s running speed and performance, ultimately setting the stages as 3.
@@ VI Conclusion and Discussion
In this paper, we have made the first attempt to explore the acquisition of high-quality images in underwater environments through 4-D LF-based imaging. To this end, we have constructed a varied and photorealistic 4-D LF-based underwater image dataset, which serves as a benchmark for future research in this area. Additionally, we presented a novel approach to enhance underwater 4-D LF images, incorporating both explicit and implicit use of depth information from 4-D LF images. Our method has demonstrated significant improvements in color correction and overall image quality through the progressive refinement of depth estimation and image enhancement. Extensive experiments have validated that our approach achieves state-of-the-art performance in enhancing 4-D LF-based underwater imaging, effectively addressing the challenges posed by light absorption and scattering in underwater environments.

### v2

@@ V-C Network Architecture Evaluation
The scores of PSNR and SSIM metrics of different methods are reported in Table II. Our proposed method demonstrates the best performance as presented. Unlike 2-D RGB-based methods, the geometric information from the 4-D LFs proves advantageous in improving underwater imaging quality. Our specific strategy enables our framework to outperform two other LF-based methods, specifically due to our meticulous design for underwater environments. As for reference time, three LF methods have the capability to enhance all views of a scene simultaneously, whereas other methods require processing each image individually.
@@ V-D Ablation Study
TABLE IV: Quantitative results of ablation study about different depth information sources, including DistgDisp, OccUnNet, and Ground Truth.
Stages PSNR $\uparrow$ SSIM $\uparrow$
DistgDisp 21.15 0.8474
OccUnNet 21.82 0.8641
Ground Truth 22.21 0.8701
TABLE V: Quantitative results of ablation study about different numbers of iterative stages. Numbers 1 to 5 represent the iterative stages
@@
2) Evaluation of Depth Estimation. Severe underwater degradation can compromise the content of images, leading to poor depth maps. As illustrated in the first and third columns of Fig. 13, the depth maps derived from estimated underwater images are subpar and lack a clear structural hierarchy. Conversely, as shown in the second and fourth columns of Fig. 13, the depth maps estimated from GT and restored image demonstrate superior results, providing sharp edges and preserving smoothness in areas of uniform depth. These results highlight the effectiveness of our framework with OccUnNet in providing better depth maps of enhanced LF images. And we compare different depth information, including ground truth and estimations from DistgDisp and OccUnNet, in Table IV. The quantitative results show that OccUnNet can achieve better performance than DistgDisp, but worse than the ground truth.
@@
We further conduct experiments on our proposed network with different stages (1, 2, 3, 4, 5). We train our model on our dataset with 5 × 5 angular resolution. Table V shows the results of different stages. Overall, adding more processing stages can enhance performance. However, beyond four stages, the improvement in enhancement performance becomes less significant. Additionally, more stages mean more parameters, which can slow down the model running. Therefore, we have made a trade-off between the model’s running speed and performance, ultimately setting the stages as 3.
V-E Limitation and Failure Cases
The currently designed algorithm does not take into account the methods for deep-sea low-light and strong noise conditions. This situation will affect the depth estimation effect, and affect the interaction of information between different perspectives, affecting the actual restoration effect. When applied to areas with diverse color changes, our method cannot restore all colors. We show one case in the supplementary materials.
@@ VI Conclusion and Discussion
In this paper, we have made the first attempt to explore the acquisition of high-quality images in underwater environments through 4-D LF-based imaging. To this end, we have constructed a varied and photorealistic 4-D LF-based underwater image dataset, which serves as a benchmark for future research in this area. Additionally, we presented a novel approach to enhance underwater 4-D LF images, incorporating both explicit and implicit use of depth information from 4-D LFs. Our method has demonstrated significant improvements in color correction and overall image quality through the progressive refinement of depth estimation and image enhancement. Extensive experiments have validated that our approach achieves state-of-the-art performance in enhancing 4-D LF-based underwater imaging, effectively addressing the challenges posed by light absorption and scattering in underwater environments.